# PVLM: Parsing-Aware Vision Language Model with Dynamic Contrastive Learning for Zero-Shot Deepfake Attribution

### v2

@@ Abstract
The challenge of tracing the source attribution of forged faces has gained significant attention due to the rapid advancement of generative models. However, existing deepfake attribution (DFA) works primarily focus on the interaction among various domains in vision modality, and other modalities such as texts and face parsing are not fully explored. Besides, they tend to fail to assess the generalization performance of deepfake attributors to unseen advanced generators like diffusion in a fine-grained manner. In this paper, we propose a novel parsing-aware vision language model with dynamic contrastive learning (PVLM) method for zero-shot deepfake attribution (ZS-DFA), which facilitates effective and fine-grained traceability to unseen advanced generators. Specifically, we conduct a novel and fine-grained ZS-DFA benchmark to evaluate the attribution performance of deepfake attributors to unseen advanced generators like diffusion. Besides, we propose an innovative parsing-guided vision language model with dynamic contrastive learning (PVLM) method to capture general and diverse attribution features. We are motivated by the observation that the preservation of source face attributes in facial images generated by GAN and diffusion models varies significantly. We employ the inherent face attributes preservation differences between GAN and diffusion face images to capture face parsing-aware forgery representations. Therefore, we devise a novel parsing encoder to focus on global face attribute embeddings, enabling parsing-guided DFA representation learning via dynamic vision-parsing matching. Additionally, we present a novel deepfake attribution contrastive center loss to pull relevant generators closer and push irrelevant ones away, which can be introduced into DFA models to enhance traceability. Experimental results show that our model exceeds the state-of-the-art on the ZS-DFA benchmark via various protocol evaluations. Codes will be available at GitHub.
@@ I Introduction
Inspired by the contrastive language-image pre-training (CLIP) model [9] which excels at zero-shot learning, we propose a parsing-aware vision language model with dynamic contrastive learning (PVLM) framework. Unlike CLIP which only focuses on image and text modals for zero-shot classification in general natural image scenarios, our PVLM model introduces fine-grained text and face parsing modalities, to enhance the learning of multiple-perspective (i.e., image, noise, and edge) visual features via cross-modality matching and dynamic cross-perspective alignment, to achieve fine-grained ZS-DFA. Specifically, our PVLM method differs from CLIP in the following points (See Fig. 2): First, since most face manipulations such as FS struggle to preserve facial attributes of the source image like eyebrows and eyes [3], as shown in Fig. 1, we employ a face parser [10] to generate parsing images to focus on different facial attribute regions. We observe that there are significant differences between pristine and forged face parsing images (See Fig. 1). Specifically, authentic face parsing images exhibit more precise attribute areas (e.g., eyes and eyebrows) than forged ones. Besides, parsing images derived from GAN-generated faces retain more accurate semantic attributes than those extracted from diffusion faces. Based on the observation, we propose a parsing encoder (PE) to capture global face parsing features, guiding our PVLM model to learn discriminative and general DFA embeddings. Second, we design a multi-perspective visual encoder (MPVE) to explore global and diverse multi-view (i.e., image, edge, and noise) forgery representations via the multi-head interaction module (MMI), as there are significant discrepancies between real and fake images in both edge and noise views [11, 12]. We present a novel dynamic cross-perspective contrastive loss to achieve dynamic vision-parsing matching. We also propose a language encoder to extract fine-grained language attribution embeddings for comprehensive vision-language matching. Finally, to enhance the ZS-DFA capability, we design a novel deepfake attribution contrastive center (DFACC) loss (See Fig. 2 (f)). Unlike the vanilla contrastive center loss [13] which enforces intra-generator compactness and inter-generator separability, our loss minimizes the distance between samples within the generator, while bringing related samples across generators closer and pushing unrelated ones farther apart. In summary, the contributions of our work are as follows:
@@
$\bullet$ We propose a novel parsing-aware vision language model with dynamic contrastive learning (PVLM) deepfake attributor, which introduces fine-grained text and face parsing modalities to enhance the learning of visual forgery features across image, noise, and edge views, thus mining both general and discriminative attribution patterns.
@@ III-C Multi-Perspective Visual Encoder
$I_{\text{v}}=I_{\text{cls}}^{\text{an}}+I_{\text{cls}}^{\text{ae}}.$ (5)
@@ III-D Parsing Encoder
To align face parsing forgery features with multi-view visual ones and adapt face parsing information to our DFA domain, we design the parsing encoder (PE) consisting of a parsing backbone with stacked convolutional layers and a parsing transformer encoder (PTE) with $E$ transformer blocks, $\text{TB}_{j}^{\text{p}}$, j=1,2,,Ej=1,2,…,E, to capture global and diverse face parsing forgery representations. As shown in Fig. 3, given $I^{\text{a}}$, we use the CelebA-HQ pre-trained face parser [10] to generate the parsing image $P\in\mathbb{R}^{3\times H\times W}$. Considering that the face parser may not accurately extract the parsing information of unseen faces, due to the domain gap between the pre-trained face dataset and the unseen one, we propose to conduct the parsing-guided multi-view DFA representation learning only during training. Specifically, PE derives the local face parsing feature $P_{\text{loc}}\in\mathbb{R}^{c\times h\times w}$ from $P$ through the backbone, to conduct the feature alignment. PE then imparts it to PTE to explore diverse long-range relations among features. In detail, $P_{\text{loc}}$ is appended with a learnable class token to capture the global parsing attribution features, to obtain $P_{\text{tok}}=\text{App}(\text{Proj}(\text{Flat}(P_{\text{loc}})))\in\mathbb{R}^{(n+1)\times d}$, where $n=\frac{chw}{chw}$ denotes the number of the parsing feature token sequence. It is then added with learnable position representations $P_{\text{e}}\in\mathbb{R}^{(n+1)\times d}$ to encode the position information, i.e.,
@@ III-E Language Encoder
To create fine and accurate text attribution prompts, we introduce the fine-grained text generator (FTG) [12]. FTG generates text prompts for each image based on hierarchical fine-grained labels provided by the GenFace dataset. To prevent information leakage from texts and considering that unseen generators may lack text prompts, we introduce fine-grained texts to train our model only during training, to conduct language-guided multi-view representation learning. Specifically, given $I^{\text{a}}$, we generate deepfake attribution prompts $T$ via FTG. To extract fine-grained language attribution-aware embeddings, we design the LE. As Fig. 3 shows, LE contains a language transformer encoder with $L$ transformer blocks, $\text{TB}_{j}^{\text{l}}$, j=1,2,,Lj=1,2,…,L. In detail, given the fine-grained deepfake attribution prompts $T$, LE first yields a series of word tokens $T^{\text{tok}}\in\mathbb{R}^{t}$ using the tokenizer [9], where $t$ represents the number of word tokens. They are then projected to the word embedding vector $T^{\text{emb}}\in\mathbb{R}^{t\times d}$, and added with learnable position embeddings $P_{\text{l}}\in\mathbb{R}^{t\times d}$, i.e., $T_{\text{1}}^{\text{tra}}={\ T}^{\text{emb}}+P_{\text{l}}$. Thereafter, they are consecutively transferred to $L$ blocks, i.e.,
@@ III-F Loss Function
$\mathcal{L}_{\text{dcpc}}=\frac{1}{b}\left(-\sum_{u=1}^{b}(y_{\text{pa}}^{u})^{T}\left[\log(S_{\text{v2v}}^{u})+\log(S_{\text{p2v}}^{u})\right]\right).$ (17)
@@
TABLE I: ZS-DFA evaluation on protocol-1. ACC scores (%) on unseen generators, after training using seen generators. ForAdapter is ForensicsAdapter [29]. Bold and underline denote the best and the second-best results.
DNA-Det 6.04 16.72 4.20 10.2 3.04 10.16 28.76 59.28 29.08 58.64 11.52 27.20 25.32 49.96 21.88 44.12 33.20 62.36 18.12 32.63
@@ IV Experiments
TABLE II: The outline of the dataset construction and protocol. # denotes the number of dataset. Pro means the protocol.
@@ IV-A Experiment Setup
Dataset collection and protocol. We collect real and fake face images created by various advanced generators, including GAN and diffusion, based on several datasets, such as GenFace [11], DF40 [30], CelebAHQ [31], and FFHQ [32]. These fake data mainly involve three manipulations, including entire face synthesis (EFS), attribute manipulation (AM), and face swap (FS). Each forgery contains some GAN-based or diffusion-based generators. We define two protocols to evaluate the attribution performance of various models under zero-shot scenarios: 1) Protocol-1 intends to investigate the performance of different attributors on unseen generators. It consists of 19 generators, such as GAN-based and diffusion-based ones from both the GenFace and DF40 datasets. We also generate the 2k fake faces using the CelebA-HQ pre-trained REFace model [33], which is a diffusion-based FS method. 2) Protocol-2 introduces additional original faces from both CelebA-HQ and FFHQ datasets based upon protocol-1.
@@
IV-B Comparison with the State of the Art
Results on protocol-1. To study the attribution performance of different models, we conduct the ZS-DFA evaluation using protocol-1. We test models on various unseen generators from DF40 after training using seen generators from GenFace. In Table I, our method exceeds most attributors by a wide gap. It also surpasses the recent network MFCLIP by about 13% ACC under the threshold of 0.7. Unlike MFCLIP which conducts the fine-grained language-guided image-noise forgery representation learning, our model introduces face parsing global features to conduct the bi-modality (e.g., text and parsing) guided multiple perspectives (e.g., image, noise, and edge) DFA representation learning. Compared to the DFA methods, the ACC of our method is nearly 4% and 13% higher than that of CPL and DNA-Det, validating the effectiveness of our model. Unlike DNA-Det, which only explores global attribution features within the image modality via patch-based contrastive learning, our model focuses on multiple modalities, including text and face parsing, as well as multiple views, such as image, noise, and edge, to achieve predominant ZS-DFA. Besides, the ACC score of models using the threshold of 0.9 is higher than that of models using the threshold of 0.7, which validates the reliability and effectiveness of the chosen threshold. We argue that a higher threshold necessitates that the model exhibit greater confidence in correct classification, thereby reducing the occurrence of misclassifications.
Results on protocol-2. To further assess the attribution performance of various methods, we conduct the ZS-DFA evaluation on protocol-2. We test models on various unseen generators and FFHQ after training using both seen generators and CelebA-HQ. In Table III, the ACC of our method exceeds that of most SOTAs, showing the superior ZS-DFA capability of our method. Specifically, the average ACC score of our PVLM model is about 19.31% higher than that of CViT on protocol-2. Unlike CViT merely explores global image forgery traces, our method focuses on global visual forgery features across three views (i.e., image, noise, and edge) and conducts fine-grained vision-language alignment and vision-parsing matching. Besides, models trained using protocol-2 outperform those trained with protocol-1. For instance, MFCLP and DNA-Det gain the average ACC score of 38.70% and 32.63% under the threshold of 0.9 setting on protocol-1, respectively. By contrast, they attain 43.38% and 50.48% ACC on protocol-2. It is noticed that a nearly 4.7% and 17.85% growth of ACC could be realized by adding original face distributions, which shows that pristine face images could facilitate the ZS-DFA performance of various deepfake attributors. We argue that models may extract more discriminative and common forgery features for each generator, due to the addition of authentic face distributions.
TABLE IV: ZS-DFA evaluation on unseen in-the-wild deepfake datasets. ACC scores (%) on unseen deepfake datasets, after training using seen generators and CelebA-HQ.
@@ IV-B Comparison with the State of the Art
Generalization to unseen in-the-wild deepfake datasets. To offer stronger evidence of domain robustness beyond synthetic benchmarks, we train networks using protocol-2 and test them on unseen in-the-wild deepfake datasets, such as WildDeepfake [37], Celeb-DF [38], and DFDC [39]. In Table IV, our model surpasses most methods across various fake face images from unseen datasets, which shows that our model achieves excellent domain robustness. To conduct the runtime benchmark, we assess the computational costs of models, including the parameter count and inference time. Although our PVLM model has a larger number of parameters than the lightweight DNA-Det, the attribution performance of PVLM is significantly improved on unseen deepfake datasets. In particular, the ACC score of PVLM is about 30% higher than that of DNA-Det, with an increase of 236.02M parameters but a reduction of 0.02s in the inference time. In the future, we aim to apply knowledge distillation and model architecture modifications, such as a window-based multi-head self-attention, to achieve faster efficiency.

### v3

# PVLM: Parsing-Aware Vision-Language Model with Dynamic Contrastive Learning for Zero-Shot Deepfake Attribution

@@ Abstract
The challenge of tracing the source attribution of forged faces has gained significant attention due to the rapid advancement of generative models. However, existing deepfake attribution (DFA) works primarily focus on the interaction among various domains in vision modality, and other modalities such as texts and face parsing are not fully explored. Besides, they tend to fail to assess the generalization performance of deepfake attributors to unseen advanced generators like diffusion in a fine-grained manner. In this paper, we propose a novel parsing-aware vision language model with dynamic contrastive learning (PVLM) method for zero-shot deepfake attribution (ZS-DFA), which facilitates effective and fine-grained traceability to unseen advanced generators. Specifically, we conduct a novel and fine-grained ZS-DFA benchmark to evaluate the attribution performance of deepfake attributors to unseen advanced generators like diffusion. Besides, we propose an innovative PVLM attributor based on the vision-language model to capture general and diverse attribution features. We are motivated by the observation that the preservation of source face attributes in facial images generated by GAN and diffusion models varies significantly. We propose to employ the inherent facial attributes preservation differences to capture face parsing-aware forgery representations. Therefore, we devise a novel parsing encoder to focus on global face attribute embeddings, enabling parsing-guided DFA representation learning via dynamic vision-parsing matching. Additionally, we present a novel deepfake attribution contrastive center loss to pull relevant generators closer and push irrelevant ones away, which can be introduced into DFA models to enhance traceability. Experimental results show that our model exceeds the state-of-the-art on the ZS-DFA benchmark via various protocol evaluations. Codes will be available at GitHub.
@@ I Introduction
Figure 3: Cross-generator correlation matrix visualization. We randomly select 10k samples from each of the two generators to calculate the FID score, to measure the relevance. The lower the FID score, the greater the correlation across generators. The darker red means stronger correlation and the lighter red denotes weaker relevance.
Figure 4: The visualization of priors from different domains including face parsing, edge, and frequency. Each column shows a face yielded by various generators. The first to forth rows represent the RGB image, face parsing image, edge image extracted by Sobel, and the frequency image derived from the fast fourier transform (FFT), respectively. We randomly select 10k real prior images and 10k fake ones to calculate the FID score. The higher the FID score, the greater the difference between the real and fake prior distribution.
@@
Inspired by the contrastive language-image pre-training (CLIP) model [26] which excels at zero-shot learning, we propose a parsing-aware vision-language model with dynamic contrastive learning (PVLM) framework. Unlike CLIP which only focuses on image and text modals for zero-shot classification in general natural image scenarios, our PVLM model introduces fine-grained text and face parsing modalities, to enhance the learning of multiple-perspective (i.e., image, noise, and edge) visual features via cross-modality matching and dynamic cross-perspective alignment, to achieve fine-grained ZS-DFA. Specifically, our PVLM method differs from CLIP in the following points (See Fig. 2): First, we design a multi-perspective visual encoder (MPVE) to explore global and diverse multi-view (i.e., image, edge, and noise) forgery representations via the multi-head interaction module (MMI), as there are significant discrepancies between real and fake images in both edge and noise views [44, 43]. Second, since most face manipulations such as FS struggle to preserve facial attributes of the source image like eyebrows and eyes [24], as shown in Fig. 1, we employ a face parser [38] to generate parsing images to focus on different facial attribute regions. We observe that there are significant differences between pristine and forged face parsing images (See Fig. 1). Specifically, authentic face parsing images exhibit more precise attribute areas (e.g., eyes and eyebrows) than forged ones. Besides, parsing images derived from GAN-generated faces retain more accurate semantic attributes than those extracted from diffusion faces. In Fig. 4, the parsing image in the diffusion-generated facial image is more discriminative than that in the GAN-synthesized one, compared to other priors extracted by edge Sobel or the frequency-based FFT. The FID score of the diffusion-generated face parsing mode is significantly higher than that of other diffusion priors, and there are evident FID score differences in face parsing images among various generators, compared to other edge and frequency distributions. This demonstrates the generalization advantage of face parsing priors over other forgery cues (e.g., frequency-domain artifacts). Based on the above observation and considering that different generators may produce various facial attribute manipulation artifacts, we propose a parsing encoder (PE) to focus on global fine-grained and diverse semantic forgery regions of facial attributes, which can adapt the face parsing information to our DFA domain and align the face parsing forged features with multi-view visual ones, guiding our PVLM model to learn discriminative and general DFA embeddings. We present a novel dynamic cross-perspective contrastive loss to achieve dynamic vision-parsing matching. We also propose a language encoder to extract fine-grained language attribution embeddings for comprehensive vision-language matching. Finally, inspired by the vanilla contrastive center loss (VCC) [25], we aim to enhance DFA representations through metric learning. However, VCC tends to enforce intra-generator compactness and inter-generator separability (see Fig. 2 (e)), which is regarded as a sub-optimal metric learning solution since there are strong correlations between some generators, with some exhibiting weak correlations (see Fig. 3), and relevant samples across generators should be closer to each other in the feature space. To address the limitation, we design a novel deepfake attribution contrastive center (DFACC) loss (See Fig. 2 (f)) to minimize the distance between samples within the generator, while bringing related samples across generators closer and pushing unrelated ones farther apart, which facilitates our model to capture generator-common artifacts to improve the generalization. In summary, the contributions of our work are as follows:
@@
$\bullet$ We propose a novel parsing-aware vision-language model with dynamic contrastive learning (PVLM) deepfake attributor, which introduces fine-grained text and face parsing modalities to enhance the learning of visual forgery features across image, noise, and edge views, thus mining both general and discriminative attribution patterns.
@@ II Related Work
Vision-Language Model. C2P-CLIP [33] presents a novel approach known as category common prompt CLIP, which injects category-specific prompts into the text encoder and enriches the image encoder with relevant category concepts via vision-language contrastive learning, resulting in enhanced detection performance. MFCLIP [43] integrates fine-grained noise forgery embeddings with global image forgery prompts, enhancing them via dynamic vision-language contrastive learning to boost generalization to face images generated by diffusion models. ForensicAdapter [3] designs an adapter to capture unique blending boundaries of forged faces, and then enhances CLIP visual tokens using a specialized interaction strategy to boost knowledge transfer between CLIP and the adapter. By contrast, based on the novel finding of face parsing feature differences between GAN and diffusion models, we employ distinct parsing features and fine-grained text prompts to enhance the visual forgery representation learning across noise, image, and edge views, to realize fine-grained ZS-DFA.
@@ III-C Multi-Perspective Visual Encoder
$I_{\text{v}}=I_{\text{cls}}^{\text{an}}\oplus I_{\text{cls}}^{\text{ae}},$ (5)
where $\oplus$ denotes the element-wise addition.
@@ III-D Parsing Encoder
To align face parsing forgery features with multi-view visual ones and adapt face parsing information to our DFA domain, we design the parsing encoder (PE) consisting of a parsing backbone with stacked convolutional layers and a parsing transformer encoder (PTE) with $E$ transformer blocks, $\text{TB}_{j}^{\text{p}}$, j=1,2,,Ej=1,2,…,E, to capture global and diverse face parsing forgery representations. As shown in Fig. 5, given $I^{\text{a}}$, we use the CelebA-HQ pre-trained face parser [38] to generate the parsing image $P\in\mathbb{R}^{3\times H\times W}$. Considering that the face parser may not accurately extract the parsing information of unseen faces, due to the domain gap between the pre-trained face dataset and the unseen one, we propose to conduct the parsing-guided multi-view DFA representation learning only during training. Specifically, PE derives the local face parsing feature $P_{\text{loc}}\in\mathbb{R}^{c\times h\times w}$ from $P$ through the backbone, to conduct the feature alignment. PE then imparts it to PTE to explore diverse long-range relations among features. In detail, $P_{\text{loc}}$ is flattened along the channel and projected to a 2D parsing token, and then appended with a learnable class token to capture the global parsing attribution features, to derive $P_{\text{tok}}=\text{App}(\text{Proj}(\text{Flat}(P_{\text{loc}})))\in\mathbb{R}^{(n+1)\times d}$, where $n=\frac{chw}{chw}$ denotes the number of the parsing token sequence and $d$ is the dimension of the parsing token. It is then added with learnable position representations $P_{\text{e}}\in\mathbb{R}^{(n+1)\times d}$ to encode the position information, i.e.,
@@ III-E Language Encoder
To create fine and accurate text attribution prompts, we introduce the fine-grained text generator (FTG) [43]. FTG generates text prompts for each image based on hierarchical fine-grained labels provided by the GenFace dataset. To prevent information leakage from texts and considering that unseen generators may lack text prompts, we introduce fine-grained texts to train our model only during training, to conduct language-guided multi-view representation learning. Specifically, given $I^{\text{a}}$, we generate deepfake attribution prompts $T$ via FTG. To extract fine-grained language attribution-aware embeddings, we design the LE. As Fig. 5 shows, LE contains a language transformer encoder with $L$ transformer blocks, $\text{TB}_{j}^{\text{l}}$, j=1,2,,Lj=1,2,…,L. In detail, given the fine-grained deepfake attribution prompts $T$, LE first yields a series of word tokens $T^{\text{tok}}\in\mathbb{R}^{t}$ using the tokenizer [26], where $t$ represents the number of word tokens. They are then projected to the word embedding vector $T^{\text{emb}}\in\mathbb{R}^{t\times d}$, and added with learnable position embeddings $P_{\text{l}}\in\mathbb{R}^{t\times d}$, i.e., $T_{\text{1}}^{\text{tra}}={\ T}^{\text{emb}}\oplus P_{\text{l}}$. Thereafter, they are consecutively transferred to $L$ blocks, i.e.,
@@ III-F Loss Function
$\mathcal{L}_{\text{v2p}}$ $=\frac{1}{b}\sum_{u=1}^{b}{-({y_{\text{pa}}^{u}})^{T}\text{log}(}S_{\text{v2p}}^{u}),$ (17)
$\mathcal{L}_{\text{p2v}}$ $=\frac{1}{b}\sum_{u=1}^{b}{-({y_{\text{pa}}^{u}})^{T}\text{log}(}S_{\text{p2v}}^{u}).$ (18)
@@ IV Experiments
TABLE I: The outline of the dataset construction and protocol. # denotes the number of dataset. Pro means the protocol.
@@
- Celeb-DF++ FR GAN LivePortrait - 1,250
- Celeb-DF++ FR Flow LIA - 1,250
- Celeb-DF++ FR GAN FSRT - 1,250
- - FR Diffusion DiffusionAct - 1,250
- Celeb-DF++ TF GAN Real3D-Portrait - 1,250
- Celeb-DF++ TF Diffusion AniTalker - 1,250
- Celeb-DF++ TF Flow FLOAT - 1,250
- Celeb-DF++ TF Flow EDTalk - 1,250
@@ IV-A Experiment Setup
Dataset collection and protocol. We collect real and fake face images created by various advanced generators, including GAN, diffusion and flow model, based on several datasets, such as GenFace [44], DF40 [39], CelebAHQ [15], Celeb-DF++ [22], and FFHQ [16]. As Table I shows, these fake data mainly involve five manipulations, including entire face synthesis (EFS), attribute manipulation (AM), face reenactment (FR), talking face (TF), and face swap (FS). Each forgery contains some GAN-based, diffusion-based, and flow-based generators. We define two protocols to evaluate the attribution performance of various models under zero-shot scenarios: 1) Protocol-1 intends to investigate the performance of different attributors on unseen generators. It consists of 28 generators, including GAN-based, diffusion-based, and flow-based ones from the GenFace, DF40, and Celeb-DF++ datasets. We generate 2k identity-swapped faces using the CelebA-HQ pre-trained REFace model [1], which is a diffusion-based FS method. We create 70k fake faces by inputting the real images from FFHQ to the FFHQ pre-trained DiffusionAct model [1], which is a diffusion-based FR method. We synthesize 2k entire manipulated faces using the flow-based FLUX model [1]. 2) Protocol-2 introduces additional original faces from both CelebA-HQ and FFHQ datasets based upon protocol-1.
@@
TABLE II: ZS-DFA evaluation on protocol-1. ACC scores (%) on unseen generators, after training using seen generators. ForAdapter is ForensicsAdapter [3]. Bold and underline denote the best and the second-best results.
DNA-Det [40] 6.04 16.72 4.20 10.2 3.04 10.16 28.76 59.28 29.08 58.64 11.52 27.20 25.32 49.96 21.88 44.12 33.20 62.36 18.12 32.63
IV-B Comparison with the State of the Art
Results on protocol-1. To study the attribution performance of different models, we conduct the ZS-DFA evaluation using protocol-1. We test models on various unseen generators from DF40 after training using seen generators from GenFace. In Table II, our method exceeds most attributors by a wide gap. It also surpasses the recent network MFCLIP by about 13% ACC under the threshold of 0.7. Unlike MFCLIP which conducts the fine-grained language-guided image-noise forgery representation learning, our model introduces face parsing global features to conduct the bi-modality (e.g., text and parsing) guided multiple perspectives (e.g., image, noise, and edge) DFA representation learning. Compared to the DFA methods, the ACC of our method is nearly 4% and 13% higher than that of CPL and DNA-Det, validating the effectiveness of our model. Unlike DNA-Det, which only explores global attribution features within the image modality via patch-based contrastive learning, our model focuses on multiple modalities, including text and face parsing, as well as multiple views, such as image, noise, and edge, to achieve predominant ZS-DFA. Besides, the ACC score of models using the threshold of 0.9 is higher than that of models using the threshold of 0.7, which validates the reliability and effectiveness of the chosen threshold. We argue that a higher threshold necessitates that the model exhibit greater confidence in correct classification, thereby reducing the occurrence of misclassifications.
Results on protocol-2. To further assess the attribution performance of various methods, we conduct the ZS-DFA evaluation on protocol-2. We test models on various unseen generators and FFHQ after training using both seen generators and CelebA-HQ. In Table III, the ACC of our method exceeds that of most SOTAs, showing the superior ZS-DFA capability of our method. Specifically, the average ACC score of our PVLM model is about 19.31% higher than that of CViT on protocol-2. Unlike CViT merely explores global image forgery traces, our method focuses on global visual forgery features across three views (i.e., image, noise, and edge) and conducts fine-grained vision-language alignment and vision-parsing matching. We also test models on various unseen generators, including LivePortrait, FSRT, LIA, DiffusionAct, AniTalker, Real3D-Portrait, FLOAT, EDTalk, and FLUX, after training using both seen generators and CelebA-HQ. In Table IV, the average ACC score of our PVLM model is about 21.08% higher than that of MFCLIP. Different from MFCLIP which merely mines global fine-grained visual forgery patterns across noise and image modalities and enhances them via vision-language matching, our PVLM method further introduces the distinct face parsing priors and captures comprehensive and complementary forged traces across image, noise, and face parsing views via vision-language alignment and flexible vision-parsing matching. Besides, models trained using protocol-2 outperform those trained with protocol-1. For instance, MFCLP and DNA-Det gain the average ACC score of 38.70% and 32.63% under the threshold of 0.9 setting on protocol-1, respectively. By contrast, they attain 43.38% and 50.48% ACC on protocol-2. It is noticed that a nearly 4.7% and 17.85% growth of ACC could be realized by adding original face distributions, which shows that pristine face images could facilitate the ZS-DFA performance of various deepfake attributors. We argue that models may extract more discriminative and common forgery features for each generator, due to the addition of authentic face distributions.
Generalization to unseen in-the-wild deepfake datasets. To offer stronger evidence of domain robustness beyond synthetic benchmarks, we train networks using protocol-2 and test them on unseen in-the-wild deepfake datasets, such as WildDeepfake [45], Celeb-DF [21], and DFDC [6]. In Table V, our model surpasses most methods across various fake face images from unseen datasets, which shows that our model achieves excellent domain robustness. To conduct the runtime benchmark, we assess the computational costs of models, including the parameter count and inference time. Although our PVLM model has a larger number of parameters than the lightweight DNA-Det, the attribution performance of PVLM is significantly improved on unseen deepfake datasets. In particular, the ACC score of PVLM is about 30% higher than that of DNA-Det, with an increase of 236.02M parameters but a reduction of 0.02s in the inference time. In the future, we aim to apply knowledge distillation and model architecture modifications, such as a window-based multi-head self-attention, to achieve faster efficiency.
@@ IV-B Comparison with the State of the Art
TABLE IV: ZS-DFA evaluation on protocol-2. ACC scores (%) on unseen generators, after training using seen generators and CelebAHQ. ForAdapter is ForensicsAdapter [3]. Bold and underline denote the best and the second-best results.
Methods LivePortrait LIA FSRT DiffusionAct EDTalk AniTalker Real3DPortrait FLOAT FLUX Average
CLIP 22.72 33.60 22.40 36.88 22.72 37.52 35.68 53.84 11.68 24.48 13.04 27.36 6.88 15.92 4.96 14.16 3.60 8.16 15.96 27.99
CViT 9.60 17.04 3.12 7.44 2.00 6.96 9.04 22.32 8.80 18.72 0.72 1.84 25.76 54.64 33.52 58.00 13.60 30.08 11.80 24.12
CAEL 21.36 43.44 31.68 60.32 29.04 58.08 8.96 21.92 27.04 54.72 25.76 51.04 21.68 44.80 5.84 15.84 6.40 15.12 19.75 40.59
MFCLIP 42.08 74.88 44.24 81.52 43.76 81.92 2.32 6.72 42.8 83.28 37.68 69.52 28.56 60.88 43.76 72.32 6.08 12.40 32.36 60.38
DNA-Det 1.92 4.24 0.00 0.32 0.16 0.64 47.52 75.36 0.00 0.02 0.00 0.08 29.36 71.44 41.28 74.56 8.64 17.76 14.32 27.16
CPL 19.56 42.71 30.79 70.54 40.02 78.95 21.04 70.85 30.41 60.79 40.13 53.79 30.46 72.88 40.37 70.84 7.20 13.97 28.89 59.48
CARZero 5.89 12.05 6.77 16.97 8.02 18.56 12.76 20.18 6.89 10.97 7.54 14.37 5.09 11.44 8.96 15.73 13.89 29.40 8.42 16.65
DE-FAKE 4.40 11.04 5.84 14.48 6.96 17.60 10.48 23.84 5.28 12.16 5.92 13.52 3.60 9.60 6.08 14.16 12.00 28.32 6.73 16.08
ForAdapter 33.44 64.16 26.00 52.88 26.08 55.44 23.76 47.20 28.96 56.32 29.44 56.56 21.92 48.32 33.52 58.88 10.08 22.48 25.91 51.36
Ours 46.88 78.08 50.40 82.56 52.08 86.88 61.76 83.60 56.24 88.40 49.12 74.88 61.84 92.32 61.92 88.48 32.00 57.92 52.47 81.46
TABLE V: ZS-DFA evaluation on unseen in-the-wild deepfake datasets. ACC scores (%) on unseen deepfake datasets, after training using seen generators and CelebA-HQ.